# DeCo-SGD: Joint Optimization of Delay Staleness and Gradient Compression Ratio for Distributed SGD

### v1

@@ 4 Mathematical Modeling Time to Accuracy in DD-EF-SGD
[…]
@@ 4.1 Time to Iteration in DD-EF-SGD
Theorem 3 (Estimation of $T_{avg}$ in DD-EF-SGD).
Assume that throughout $t$ iterations where $\tau\ll t\ll\mathbf{T}$, the constants $a$, $b$, $\delta$, and $\tau$ remain fixed. Then the average iteration time $T_{avg}$ satisfies the following approximation:
where the approximation error is upper bounded by $O(\frac{1}{t})$.
Remark 4 (The optimal compression ratio $\delta^{*}$).
We can minimize $T_{avg}$ to $T_{comp}$ by enlarging $\tau$ and choosing a small $\delta$ Let $\delta^{*}(\tau)=\min\left\{(\tau T_{comp}-b)\frac{a}{S_{g}},T_{comp}\frac{a}{S_{g}},1\right\}$, we can obtain that $\delta\leq\delta^{*}(\tau)$. Given $\tau\geq\frac{b}{T_{comp}}$, $\delta^{*}=\delta^{*}(\tau)$ This is due to the fact that the more aggressive $\delta$ cannot get the performance gain, which is covered by the delay while leading to accuracy degradation. In this way, $\delta^{*}$ is the maximal value in the given range.
@@ 4.2 DeCo: Joint Optimization for Delay Staleness and Gradient Compression in DML
Input: $S_{g}$, $a$, $b$, $T_{comp}$
@@
for $\tau\in\{\lceil\frac{b+S_{g}/a}{T_{comp}}\rceil,\cdots,1+\lceil\frac{b}{T_{comp}}\rceil,\lceil\frac{b}{T_{comp}}\rceil\}$  do
      $\delta=\min\left\{(\tau T_{comp}-b)\frac{a}{S_{g}},T_{comp}\frac{a}{S_{g}},1\right\}$;
     /* Calculate the value of $\phi$/
@@
Input: $n$, $\gamma$, $S_{g}$, $T_{comp}$, initial parameters $\textbf{x}_{0}$, initial local error $\textbf{e}_{0}^{i}=\textbf{0}_{d}$, the frequency to update DeCo $E$
Output: $\textbf{x}_{T}$
for $i=1,\ldots,n$ do
       if $t\mod E==1$ then
             /* Update $\tau$ and $\delta$ with DeCo per $E$ iterations */
             Get $a$, $b$ from the network;
             $\tau$, $\delta$ = DeCo($S_{g}$, $a$, $b$, $T_{comp}$);
       end if
      $\hat{\Delta}_{t}^{i}=\textbf{C}_{\delta}(\textbf{e}_{t-\tau}^{i}+\textbf{g}_{t-\tau}^{i})$;
       $\textbf{e}_{t+1-\tau}^{i}=\textbf{e}_{t-\tau}^{i}+\textbf{g}_{t-\tau}^{i}-\hat{\Delta}_{t}^{i}$;
       $\textbf{x}_{t+1}=\textbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\hat{\Delta}_{t}^{i}$;
end for
Return $\textbf{x}_{T}$;
Algorithm 2 DeCo-SGD
Our design goal is to get the minimal end-to-end time to achieve a specific accuracy, which is equivalent to solving the minimum of the binary function $\phi$. Since $\tau$ is discrete, the minimum of $\phi$ cannot be directly solved by the partial derivatives. We use the critical condition $T_{avg}=T_{comp}$ (i.e., there are no redundant bubbles in the pipeline) to reduce the global optima into a countable number of local optima. The mathematical description of this goal is
$\min\ \phi(\delta,\tau)=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}},\quad\textit{s.t.}\ T_{avg}=T_{comp},\quad\tau\in\mathbb{N},\quad\delta\in\left(0,1\right].$ (10)
@@
$\min\ \phi(\delta^{*}(\tau),\tau)=\frac{1-\delta^{*}(\tau)}{\delta^{*}(\tau)(1-\frac{\delta^{*}(\tau)}{2})^{\tau}}\quad\textit{s.t.}\ \tau\in\left[\lceil\frac{b}{T_{comp}}\rceil,\cdots,\lceil\frac{b+\frac{S_{g}}{a}}{T_{comp}}\rceil\right].$ (11)
We traverse the range of $\tau$ to get $\phi^{*}$ to achieve the fastest convergence rate and choose the smallest $\tau$ as $\tau^{*}$ The procedure for obtaining $\tau^{*}$ and $\delta^{*}$ is referred to as the DeCo algorithm (shown in Algo. 1).
The implementation of DeCo in D-SGD is called DeCo-SGD (pseudo-code shown in Algo. 2 and design shown in Fig. 3). DeCo-SGD sets a frequency $E$ to update DeCo. $E$ describes the sensitivity of DeCo. To enable frequent adjustments of the compression ratio and delay staleness in response to a dynamic network environment, a smaller $E$ value can be adopted, such as $E=1$. In the normal case, a middle $E$ like $100$ is enough. DeCo-SGD incurs a negligible extra cost compared to vanilla DD-EF-SGD and the extra time complexity is $\mathcal{O}(\frac{T}{E})$. The increase in $n$ does not result in additional cost.
@@ 5.1 Experimental Settings
We evaluate our approach on diverse CV and NLP tasks using CNN lecun1998gradient , ViT vit , and GPT gpt2 models, including FashionMNIST xiao2017fashion , CIFAR-$10$, ImageNet imagenet , and Wikitext wikitext . Our baselines include D-SGD, Accordion accordion , DGA DGA , and CocktailSGD CocktailSGD . CocktailSGD is a SOTA static strategy that combines gradient compression and delayed aggregation using fixed parameters. If not specified, all compressors are Top-$k$ gradient compression dagctmc , except for CocktailSGD.
@@ 5.2 Comparison of Training Speed
As illustrated in Fig. 4, DeCo-SGD demonstrates significant improvements in training speed compared to D-SGD, Accordion, DGA, and CocktailSGD across all tasks. All experiments were conducted with $4$ worker nodes. Take ViT@ImageNet as an example, when the accuracy reaches 70%, DeCo-SGD achieves a 4.99×4.$9$ speedup444speed-up is defined as the ratio of the baseline method’s training time to that of DeCo-SGD. over D-SGD and a 1.34×1.34$\times$1.34 × speedup over CocktailSGD. At 80% accuracy, the speedup increases to 5.01×5.01$\times$5.01 × and 1.37×1.37$\times$1.37 ×, respectively. Compared with other baseline methods, the benefits of DeCo-SGD are particularly obvious, mainly for two reasons: (1) their parameter settings are suboptimal; (2) they cannot adapt to the dynamic network environment in real time. In contrast, DeCo-SGD utilizes theoretical guidance and runtime feedback to jointly optimize compression and latency, thereby achieving a more effective communication-computing trade-off throughout the training process.
Figure 4: Training time comparison across different model-dataset pairs.
@@ 5.3 Scalability Verification of DeCo-SGD
DeCo-SGD is inherently scalable, as the computational complexity of DeCo-SGD in finding $\delta^{*}$ and $\tau^{*}$ is theoretically independent of the number of worker nodes. Fig. 5 illustrates the performance of DeCo-SGD across different numbers of worker nodes, scaling from $n=4$ to $n=32$. All experiments were conducted under a fixed latency of $200$ms, with bandwidth fluctuating around $100$Mbps.
The results demonstrate that DeCo-SGD consistently maintains performance improvements as the number of nodes increases. Specifically, for GPT@Wikitext, DeCo-SGD achieves up to 3.77×3.$7$ and 1.21×1.21$\times$1.21 × speed-up over D-SGD and CocktailSGD, respectively, even at $n=32$. Similarly, for ViT@ImageNet, DeCo-SGD achieves up to 5.07×5.07$\times$5.07 × speed-up over D-SGD and 1.37×1.37$\times$1.37 × over CocktailSGD. These results highlight the adaptability of DeCo-SGD to large-scale and high-precision distributed training tasks.
@@ 5.4 DeCo-SGD under Different Bandwidths and Latency
Table 1 presents the training times of different methods on GPT@Wikitext and ViT@ImageNet under varying network conditions, with all experiments conducted using $4$ worker nodes. The numbers in the table denote the training time (in seconds) required by each method to achieve a target perplexity ($22$ for the GPT model) or accuracy ($80$% for the ViT model). The results clearly demonstrate that under all network conditions, DeCo-SGD consistently outperforms other methods, achieving the lowest training times. This highlights its robustness against bandwidth and latency variations as well as its superior communication efficiency in challenging network environments. For instance, DeCo-SGD achieves up to 5.02×5.02$\times$5.02 × and 1.30×1.30$\times$1.30 × speed-ups over D-SGD and CocktailSGD, respectively.
@@
Table 1: Comparison of training time (s) under different methods and network conditions. $a$ is the average bandwidth and $b$ is the end-to-end latency. Detailed numbers are in the Appendix.
Model @DataSet a(Gbps), b(s) D-SGD Accordion DGA CocktailSGD DeCo-SGD
GPT @Wikitext $0.1,0.1$ $6396.95\pm 32.36$ $1955.28\pm 13.59$ $5506.78\pm 26.77$ $1525.21\pm 12.91$ $\mathbf{1306.29}\pm\mathbf{10.29}$
$0.5,0.1$ $1571.37\pm 11.13$ $829.06\pm 6.65$ $1346.34\pm 10.75$ $712.93\pm 5.38$ $\mathbf{642.79}\pm\mathbf{5.74}$
$0.1,1.0$ $7232.88\pm 35.62$ $2510.26\pm 22.73$ $5764.60\pm 29.69$ $1683.11\pm 14.66$ $\mathbf{1440.59}\pm\mathbf{11.87}$
$0.5,1.0$ $2363.73\pm 13.87$ $1325.12\pm 10.90$ $1562.93\pm 11.89$ $1132.63\pm 9.91$ $\mathbf{997.25}\pm\mathbf{7.92}$
ViT @ImageNet $0.1,0.1$ $1373.84\pm 10.58$ $440.14\pm 3.28$ $1183.68\pm 7.85$ $364.77\pm 1.02$ $\mathbf{283.06}\pm\mathbf{1.80}$
$0.5,0.1$ $446.20\pm 3.39$ $274.60\pm 2.21$ $385.60\pm 1.57$ $219.39\pm 0.98$ $\mathbf{169.13}\pm\mathbf{0.47}$
$0.1,1.0$ $1530.57\pm 9.89$ $604.67\pm 2.79$ $1287.03\pm 10.70$ $505.91\pm 2.56$ $\mathbf{397.16}\pm\mathbf{1.39}$
$0.5,1.0$ $631.15\pm 3.33$ $373.79\pm 1.92$ $444.38\pm 2.18$ $278.57\pm 1.35$ $\mathbf{235.95}\pm\mathbf{0.88}$

### v2

@@ 4 Mathematical Modeling Time to Accuracy in DD-EF-SGD
We denote the bandwidth (bits/s) as a time-varying function $a(s)$, where $s$ represents continuous time, the average bandwidth over the previous $E$ iterations as $\bar{a}$, end-to-end latency (s) as $b$, and gradient size (bits) as $V_{g}$. We denote the computation time per iteration (s) and the average end-to-end training time per iteration (s) as $T_{\text{comp}}$ and $T_{\text{avg}}$, respectively. We use $\lceil\cdot\rceil$ to represent the ceiling function. Detailed proofs are provided in the Appendix.
4.1 Iteration Timing Model of DD-EF-SGD
In DD-EF-SGD with one worker and server, each iteration consists of three stages: local computation $T_{\text{comp}}$; gradient transmission time $t_{\delta}$, where $\int_{0}^{t_{\delta}}a(s)\,\mathrm{d}s{=}\delta V_{g}$ and fixed end-to-end latency $b$.
Let $t$ denote the iteration index ($t=0,1,2,\dots$). Considering the timing constraint under delayed aggregation DGA, the (t+1)(t{+}1)-th computation starts after both the $t$-th computation and the $(t{-}\tau)$-th update are completed. Let $TS_{t}$ and $TC_{t}$ be the completion times of the $t$-th computation and communication, respectively. This leads to the following recursive relation:
$TS_{t+1}=T_{\text{comp}}+\max\{\,TS_{t},\;TC_{t-\tau}\,\}.$ (8)
4.2 Modeling Assumptions
To facilitate the time-to-iteration model in DD-EF-SGD  we introduce the following assumptions.
Assumption 5 (Bottleneck equivalence). In synchronous parameter-server systems, the overall iteration time is dominated by the slowest link. This allows us to analyze the multi-worker system by focusing on a single worker-server link representing the bottleneck, without loss of generality for the analysis.
Assumption 6 (Historical average bandwidth estimation). Based on network measurement studies strauss2003measurement; li2004first, network bandwidth exhibits local stationarity at minute-level timescales, with average bandwidth prediction errors typically within $10\%$ ng2002predicting. We define the window for recording bandwidth as $E$ iterations. Therefore, we let $a(s)=\bar{a}$ in the following $E$ iterations, Every $E$ iterations, we update the bandwidth. That means the gradient transmission time $t_{\delta}{=}\frac{\delta V_{g}}{\bar{a}}$.
@@ 4.3 Time to Iteration in DD-EF-SGD
Theorem 3 (Estimation of $T_{\text{avg}}$ in DD-EF-SGD).
Under Assumptions 5, 6, the average iteration time $T_{\text{avg}}$ satisfies the following approximation:
$T_{\text{avg}}\approx\max{\left\{\frac{T_{\text{comp}}+b+\delta\cdot V_{g}/\bar{a}}{\tau+1},\frac{\delta\cdot V_{g}}{\bar{a}},T_{\text{comp}}\right\}},$ (9)
where the approximation error is caused by the pipeline boundary and can be neglected when $E\gg\tau$.
Remark 4 (The locally optimal compression ratio $\delta^{*}(\tau)$).
Our goal is to minimize the iteration time $T_{\text{avg}}$ without strictly sacrificing model accuracy. According to Eq. (9), the lower bound of the average iteration time is the computation time $T_{\text{comp}}$. There exists a locally optimal compression ratio given a fixed $\tau$, denoted as $\delta^{*}(\tau)$, such that $T_{\text{avg}}(\tau,\delta^{*}(\tau))=T_{\text{comp}}$. Choosing any $\delta\leq\delta^{*}$ yields no gain in $T_{\text{avg}}$, but degrade the model performance due to more information loss. Thus, for a $\tau$, $\delta^{*}(\tau)=\min\left\{\max\left(0,(\tau T_{\text{comp}}-b)\frac{\bar{a}}{V_{g}}\right),T_{\text{comp}}\frac{\bar{a}}{V_{g}},1\right\}$.
@@ 4.4 DeCo: Joint Optimization for Delay Staleness and Gradient Compression in DML
Input: $n$, $\gamma$, $\mathbf{T}$, $C(\cdot)$, $V_{g}$, $T_{\text{comp}}$, $E$
Output: $\textbf{x}_{\mathbf{T}}$
Initialize $\mathbf{x}_{0}$, $\mathbf{e}_{0}^{i}=\mathbf{0}_{d}$;
Initialize $\tau$ and $\delta$ based on DeCo algorithm;
for $t\in[\mathbf{T}]$ do
 /* Worker side */
 for $i\in[n]$ do
      Worker $i$ waits until it receives $\textbf{x}_{t}$;
      Calculate $\mathbf{g}_{t}^{i}$ based on the global model $\mathbf{x}_{t}$;
    $\Delta_{t}^{i}=\textbf{C}_{\delta}(\textbf{e}_{t}^{i}+\gamma\textbf{g}_{t}^{i})$;
    $\textbf{e}_{t+1}^{i}=\textbf{e}_{t}^{i}+\gamma\textbf{g}_{t}^{i}-\Delta_{t}^{i}$;
      Send $\Delta_{t}^{i}$ to the server;
   end for
 /* Server side */
   Server waits until it receives $\Delta_{t-\tau}^{i}$ for all $i\in[n]$;
 $\textbf{x}_{t+1}=\textbf{x}_{t}-\frac{1}{n}\sum_{i=1}^{n}\Delta_{t-\tau}^{i}$;
   Broadcast $\textbf{x}_{t+1}$;
 /* Update $\tau$ and $\delta$ with DeCo */
 if $(t+1)\mod E==0$ then
      Get $\bar{a}$, $b$ from the network;
    $\tau$, $\delta$ = DeCo($V_{g}$, $\bar{a}$, $b$, $T_{\text{comp}}$);
      Broadcast new $\tau$ and $\delta$ to workers;
   end if
end for
Return $\textbf{x}_{\mathbf{T}}$;
Algorithm 2 DeCo-SGD
Input: $V_{g}$, $\bar{a}$, $b$, $T_{\text{comp}}$
@@
for $\tau\in\{\lceil\frac{b+V_{g}/\bar{a}}{T_{\text{comp}}}\rceil,\cdots,1+\lceil\frac{b}{T_{\text{comp}}}\rceil,\lceil\frac{b}{T_{\text{comp}}}\rceil\}$ do
   $\delta=\min\left\{(\tau T_{\text{comp}}-b)\frac{\bar{a}}{V_{g}},T_{\text{comp}}\frac{\bar{a}}{V_{g}},1\right\}$;
  /* Calculate the value of $\phi$ */
@@
Our design goal is to get the minimal end-to-end time to achieve a specific accuracy, which is equivalent to solving the minimum of the binary function $\phi$. Since $\tau$ is discrete, the minimum of $\phi$ cannot be directly solved by the partial derivatives. We use the critical condition $T_{\text{avg}}=T_{comp}$ (i.e., there are no redundant bubbles in the pipeline) to reduce the global optima into a countable number of local optima. The mathematical description of this goal is
$\min\ \phi(\delta,\tau)=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}},\textit{s.t.}\ T_{\text{avg}}{=}T_{\text{comp}},\tau\in\mathbb{N},\delta\in\left(0,1\right].$ (10)
@@
$\min\ \phi(\delta^{*}(\tau),\tau)=\frac{1-\delta^{*}(\tau)}{\delta^{*}(\tau)(1-\frac{\delta^{*}(\tau)}{2})^{\tau}}$ (11)
$\textit{s.t.}\ \tau\in\left[\lceil\frac{b}{T_{comp}}\rceil,\cdots,\lceil\frac{b+\frac{V_{g}}{\bar{a}}}{T_{comp}}\rceil\right].$
We traverse the range of $\tau$ to get $\phi^{*}$ to achieve the fastest convergence rate and choose the smallest $\tau$ as $\tau^{*}$ . The procedure for obtaining $\tau^{*}$ and $\delta^{*}$ is referred to as the DeCo algorithm (shown in Algo. 3).
Figure 5: System implementation of DeCo-SGD. Each worker has: (1) Local compute unit to local compute and compress gradients; (2) Network monitor to get the bandwidth and latency; The server has a cache to save staled updates, the compute unit to update the model based on staled updates and DeCo Unit to decide ($\delta,\tau$) based on the network conditions. Workers and the server communicate with each other with Comm. API.
4.5 Efficient Implementation of DeCo
The implementation of DeCo in D-SGD is integrated into DeCo-SGD (pseudo-code shown in Algo. 2, design in Fig. 4, and system implementation in Fig. 5). We use a pre-computed Look-up Table (LUT) strategy to ensure its decision-making process runs in $\mathcal{O}(1)$ runtime complexity. The implementation consists of two operational phases:
4.5.1 Pre-computation
To build the LUT, DeCo-SGD profiles the network during a warm-up phase to estimate the bandwidth range $a_{min},a_{max}]$. This range is then linearly interpolated into several sample points (like $20$ points). With measured $b$, the server runs DeCo offline for each sample point to compute and store the optimal configuration.
4.5.2 Runtime Lookup
During training, the worker updates the average bandwidth $\bar{a}$ per $E$ iterations. The system then queries the LUT for the optimal configuration. If $\bar{a}$ is within the recorded range, the system retrieves $(\delta^{*},\tau^{*})$ by rounding $\bar{a}$ to the nearest discretized bandwidth in the LUT. This $O(1)$ lookup ensures minimal runtime cost. If $\bar{a}$ falls outside the recorded range $a_{\min},a_{\max}]$, DeCo-SGD falls back to online computation in the server to ensure stability and correctness.
@@ 5 Evaluation Experiments
5.1 Experimental Environment
The experiments are conducted on a distributed cluster comprising $32$ nodes. Each node is equipped with an Ubuntu 24.04 LTS system, an Intel Xeon Gold $6230$ CPU, and a single Nvidia A4040 GPU with 4848GB memory. The Python version is $3.10.16$, and other used libraries are all based on the Python version. We use Python $3.10.16$, PyTorch $2.5.1$ with CUDA $12.4$ as the ML toolkit and Gloo as the communication backend.
@@ 5.2 Experimental Settings
Experiment tasks: To demonstrate the superiority of our design, we evaluate three representative tasks: CIFAR-$10$ with ResNet-34 resnet (denoted as ResNet@CIFAR-$10$), ImageNet imagenet using ViT vit (denoted as ViT@ImageNet) and Wikitext wikitext on GPT gpt2 (denoted as GPT@Wikitext). The ResNet architecture is ResNet-34, used widely in previous works dagctmc; stc. The ViT base model features approximately $86$ million parameters, while the smallest GPT variant contains $124$ million parameters. For ResNet@CIFAR-$10$, the learning rate is $0.01$, and the batch size per node is $32$. For ViT@ImageNet and GPT@Wikitext, the learning rates are $0.01$ and $0.1$, respectively, with batch sizes of $40$ and $5$. Regarding the initialization strategy, we train ResNet from scratch, whereas ViT and GPT are initialized with pre-trained weights.
Baselines: We compare D-SGD, Accordion, DAGC, DC2, DGA, CocktailSGD as baselines. D-SGD is the case $\tau=0,\delta=1$ . DAGC dagctmc , Accordion accordion, DC2 dc2 are related adaptive gradient compression algorithms. DAGC is the SOTA compressor in communication-constrained non-IID scenarios, assigning different compression ratios to workers with different local dataset volumes. $\bar{\delta}$ of DAGC is equal to intial $\delta$ of DeCo-SGD. Accordion is the SOTA adaptive gradient compression algorithm, which sets small $\delta$ in the critical iterations and sets large $\delta$ if not. The large $\delta$ (as well as small $\delta$) is equal to the $\delta$ in DeCo-SGD under the constraint bandwidth (the half of the constraint bandwidth). DC2 adapts $\delta$ based on the network condition. We use DA2 in DC2 and $k_{0}$ is equal to intial $\delta$ of DeCo-SGD. The compressor we use in all baselines is Top-$k$, which transmits the largest $k$ elements of gradients. We select Top-$k$ for its empirical robustness and widespread adoption in DML. DGA first introduces delayed aggregation in FL to reduce the negative effect of the high latency. We set K=1K{=}1 in DGA to close the infrequent communication and $D$ equal to the initial $\tau$ of DeCo-SGD. CocktailSGD wang2023cocktailsgd uses the fixed $\tau{=}1$ and the compressor composed by Random-$k$ with $k{=}20\%$, Top-$k$ with $k{=}10\%$ and using $4$-bit quantization, as stated in the main body.
Figure 6: The network bandwidth over time with an constraint of 100100Mbps.
Figure 7: Training time comparison across different tasks.
Network environment: We simulate challenging network environments (static high latency $\geq$ 100100ms, dynamic low bandwidth, constrained below 11Gbps). We measure the performance of our design under mulitple network conditions in Section $5.$ An example of dynamic bandwidth over time is shown in Fig. 6.
@@ 5.3 Comparison of Training Speed
As illustrated in Fig. 7, DeCo-SGD demonstrates significant improvements in training speed compared to D-SGD, Accordion, DGA, and CocktailSGD in all tasks. All experiments were conducted with $4$ worker nodes, under a fixed latency of 200200ms, with bandwidth fluctuating around 100100Mbps. Taking ViT@ImageNet as an example, when accuracy reaches 70%, DeCo-SGD achieves a $4.99\times$ speedup [^3] Speed-up is defined as the ratio of the baseline method’s training time to that of DeCo-SGD. over D-SGD and a $1.34\times$ speedup over CocktailSGD. At 80% accuracy, the speedup increases to $5.01\times$ and $1.37\times$, respectively. Compared with other baseline methods, the benefits of DeCo-SGD are particularly obvious, mainly for two reasons: (1) their parameter settings are suboptimal; (2) they cannot adapt to the dynamic network environment in real time. In contrast, DeCo-SGD utilizes theoretical guidance and runtime feedback to jointly optimize compression and latency, thereby achieving a more effective communication-computing trade-off throughout the training process.
@@ 5.4 Scalability Verification of DeCo-SGD
DeCo-SGD is inherently scalable, as the computational complexity of DeCo-SGD in finding $\delta^{*}$ and $\tau^{*}$ is theoretically independent of the number of worker nodes. Fig. 8 illustrates the performance of DeCo-SGD across different numbers of worker nodes, scaling from n=4n{=}4 to n=32n{=}32. The results demonstrate that DeCo-SGD consistently maintains performance improvements as the number of nodes increases. Specifically, for GPT@Wikitext, DeCo-SGD achieves up to $3.77\times$ and $1.21\times$ speed-up over D-SGD and CocktailSGD, respectively, even at n=32n{=}32. Similarly, for ViT@ImageNet, DeCo-SGD achieves up to $5.07\times$ speed-up over D-SGD and $1.37\times$ over CocktailSGD.
@@ 5.5 DeCo-SGD under Different Bandwidths and Latency
Table 3 presents the training times of different methods on GPT@Wikitext and ViT@ImageNet under varying network conditions, with all experiments conducted using $4$ worker nodes, using D-SGD, Accordion, DC2, DGA, CocktialSGD as baselines. Due to space limitations, we report the training time to reach a representative target metric for each task: $22$ for GPT and 8080% accuracy for ViT. Detailed convergence curves covering the entire training process are provided in the Appendix, which verify that DeCo-SGD consistently outperforms other methods, not just the specific points reported here. The advantage increases further as bandwidth decreases and network conditions worsen (bandwidth drops, latency increases). The results clearly demonstrate the robustness of our design against bandwidth and latency variations. The advantage of our design over baselines increases as network conditions worsen. We achieve up to $5.02\times$ and $1.30\times$ speed-ups over D-SGD and CocktailSGD, respectively. Notablely, in dynamic network environments, DC2 performs better than Accordion, but it does not address the problem of high latency. This shows that gradient compression alone cannot solve the problem.
@@
Table 3: Comparison of training time (s) under different methods and network conditions. $a$ is maximum bandwidth and $b$ is the end-to-end latency. The numbers in parentheses represent the speedup compared to DeCo-SGD.
Model
@DataSet
a(Gbps),
b(s)
D-SGD Accordion DC2 DGA CocktailSGD DeCo-SGD
GPT
@Wikitext
$0.1,0.1$ $6396.95\pm 32.36$ ($4.90\times$) $1955.28\pm 13.59$ ($1.50\times$) $1901.13\pm 11.45$ ($1.45\times$) $5506.78\pm 26.77$ ($4.22\times$) $1535.21\pm 12.91$ ($1.18\times$) $\mathbf{1306.29\pm 10.29}$
$0.5,0.1$ $1571.37\pm 11.13$ ($2.44\times$) $829.06\pm 6.65$ ($1.29\times$) $801.72\pm 6.06$ ($1.24\times$) $1346.34\pm 10.75$ ($2.09\times$) $712.93\pm 5.38$ ($1.11\times$) $\mathbf{642.79\pm 5.74}$
$0.1,1.0$ $7232.88\pm 35.62$ ($5.02\times$) $2510.26\pm 22.73$ ($1.74\times$) $2469.74\pm 22.98$ ($1.71\times$) $5764.60\pm 29.69$ ($4.00\times$) $1683.11\pm 14.66$ ($1.17\times$) $\mathbf{1440.59\pm 11.87}$
$0.5,1.0$ $2363.73\pm 13.87$ ($2.37\times$) $1325.12\pm 10.90$ ($1.33\times$) $1321.25\pm 10.31$ ($1.32\times$) $1562.93\pm 11.89$ ($1.57\times$) $1132.63\pm 9.91$ ($1.14\times$) $\mathbf{997.25\pm 7.92}$
ViT
@ImageNet
$0.1,0.1$ $1373.84\pm 10.58$ ($4.85\times$) $440.14\pm 3.28$ ($1.55\times$) $415.69\pm 3.14$ ($1.46\times$) $1183.68\pm 7.85$ ($4.18\times$) $364.77\pm 1.02$ ($1.29\times$) $\mathbf{283.06\pm 1.80}$
$0.5,0.1$ $446.20\pm 3.39$ ($2.64\times$) $274.60\pm 2.21$ ($1.62\times$) $253.36\pm 2.17$ ($1.49\times$) $385.60\pm 1.57$ ($2.28\times$) $219.39\pm 0.98$ ($1.30\times$) $\mathbf{169.13\pm 0.47}$
$0.1,1.0$ $1530.57\pm 9.89$ ($3.85\times$) $604.67\pm 2.79$ ($1.52\times$) $579.03\pm 2.53$ ($1.45\times$) $1287.03\pm 10.70$ ($3.24\times$) $505.91\pm 2.56$ ($1.27\times$) $\mathbf{397.16\pm 1.39}$
$0.5,1.0$ $631.15\pm 3.33$ ($2.67\times$) $373.79\pm 1.92$ ($1.58\times$) $368.29\pm 1.83$ ($1.56\times$) $444.38\pm 2.18$ ($1.88\times$) $278.57\pm 1.35$ ($1.18\times$) $\mathbf{235.95\pm 0.88}$
Table 4: Comparison of training time (s) under different methods and network conditions in non-IID scenarios.
Model
@DataSet
a(Gbps),
b(s)
D-SGD DAGC DC2 DGA CocktailSGD DeCo-SGD
GPT
@Wikitext
$0.1,0.1$ $7276.34\pm 49.33$ ($5.13\times$) $2283.53\pm 12.55$ ($1.61\times$) $2244.73\pm 12.21$ ($1.58\times$) $6108.13\pm 35.23$ ($4.31\times$) $1732.25\pm 12.58$ ($1.22\times$) $\mathbf{1417.55\pm 11.23}$
$0.5,0.1$ $1785.64\pm 12.55$ ($2.41\times$) $969.11\pm 7.12$ ($1.30\times$) $917.54\pm 7.42$ ($1.23\times$) $1515.61\pm 12.48$ ($2.04\times$) $815.52\pm 6.45$ ($1.10\times$) $\mathbf{741.35\pm 5.73}$
$0.1,1.0$ $8179.46\pm 54.32$ ($4.76\times$) $2841.85\pm 20.07$ ($1.65\times$) $2827.36\pm 20.91$ ($1.64\times$) $6517.52\pm 34.66$ ($3.79\times$) $2018.44\pm 14.65$ ($1.17\times$) $\mathbf{1718.41\pm 13.46}$
$0.5,1.0$ $2736.48\pm 19.64$ ($2.31\times$) $1576.29\pm 12.98$ ($1.32\times$) $1570.48\pm 12.07$ ($1.32\times$) $1775.52\pm 13.29$ ($1.50\times$) $1459.16\pm 11.63$ ($1.23\times$) $\mathbf{1186.33\pm 9.27}$
ViT
@ImageNet
$0.1,0.1$ $1273.93\pm 9.32$ ($4.64\times$) $567.73\pm 3.73$ ($2.06\times$) $537.85\pm 3.14$ ($1.95\times$) $1324.88\pm 10.26$ ($4.82\times$) $368.39\pm 1.82$ ($1.34\times$) $\mathbf{274.76\pm 1.45}$
$0.5,0.1$ $434.16\pm 2.15$ ($2.34\times$) $364.22\pm 1.80$ ($1.93\times$) $353.70\pm 1.79$ ($1.91\times$) $467.97\pm 2.10$ ($2.52\times$) $224.09\pm 1.38$ ($1.21\times$) $\mathbf{185.90\pm 0.68}$
$0.1,1.0$ $1724.74\pm 10.47$ ($4.22\times$) $672.44\pm 3.21$ ($1.64\times$) $644.02\pm 3.14$ ($1.57\times$) $1367.39\pm 8.25$ ($3.34\times$) $555.26\pm 3.11$ ($1.36\times$) $\mathbf{409.18\pm 2.48}$
$0.5,1.0$ $677.23\pm 4.21$ ($2.78\times$) $489.41\pm 2.21$ ($2.00\times$) $478.93\pm 1.85$ ($1.96\times$) $599.93\pm 2.83$ ($2.46\times$) $283.89\pm 1.62$ ($1.17\times$) $\mathbf{243.52\pm 1.02}$
5.6 DeCo-SGD under non-IID Scenarios
Federated learning or other data silo scenarios encounter non-IID problems, leading to accuracy degradation niid2020icml. We compare DeCo-SGD with other algorithms in non-IID settings. To evaluate performance under these conditions, we simulate the label imbalance and quantity imbalance using a Dirichlet distribution with a concentration parameter of $0.5$ niidbench. Notably, for these experiments, we replace the Accordion with DAGC compared to Section 5.5, a more competitive baseline specifically designed for non-IID scenarios. In non-IID scenarios, we replace $\phi$ with $\phi^{\prime}=\phi/\delta$ according to Remark 1. Despite the increased training uncertainty and instability inherent in non-IID data, DeCo-SGD demonstrates exceptional adaptability. It achieves speed-ups ranging from $1.23\times$ to $2.06\times$ over the strong baseline DAGC, significantly outperforming the gains observed in IID settings. In non-IID scenarios, our design achieves up to $5.13\times$ and $1.36\times$ speed-ups over D-SGD and CocktailSGD.